OpenStudy (anonymous):

5sqrtx-7=10 x = -5 x = -3 x = 9 x = 11

OpenStudy (anonymous):

umm none of those. is it written right? \[5\sqrt{x}-7=10\] x would be 289/25

OpenStudy (anonymous):

also we know it cant be a negative number since the negative under the radical would be imaginary

OpenStudy (anonymous):

oh unless x-7 is all under the radical

OpenStudy (anonymous):

yes it is wriitten rite

OpenStudy (anonymous):

its x-7=10

OpenStudy (anonymous):

i see, use parenthesis to clarify next time \[\sqrt{x-7}=(10/5)\] the 10/5 reduces to 2, square each side to remove the square root so: x-7=4 then x=4+7=11

OpenStudy (anonymous):

so x=11

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank u